# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v7

@@ 3.1 Conics and the projective plane
where the superscript T denotes the transpose, therefore these two associated line-point correlations coincide exactly when the point-line correlation is represented by a symmetric matrix $A=A^{T}$ which is known as a polarity. A polarity can also be characterized as a correlation that is also an involution.
@@ 3.1.1 Pencils of conics
Assume $A$ is regular. The pencil, with the exception of $A$, can be written $A+tB,t\in\mathbb{R}$. $A+tB$ is a degenerate conic exactly when $|A+tB|=0$. The left-hand side is a cubic polynomial in $t$, hence has at most 3 roots. ∎
@@ 3.1.2 Double contact
when $P$ is incident to $p$. In particular, the intersection points of $B$ with $A$ are always the intersection points of $p$ with $A$ and the tangent lines there are always the same, independently of $c$. Thus, $A$ and $B$ are said to be in double contact with the chord of contact of the line of contact $p$, and the pencil of conics defined by Eq. (6) is called a double contact pencil. The property of being in double contact is clearly self-dual. Indeed, two point-wise conics share tangent lines at two common points if and only if the two corresponding line-wise conics share turning points at two common lines. Furthermore, Eq. (6) is self-dual, that is, the line-wise conics of a double contact pencil can be expressed as the linear span of a regular conic and a double point, given by the meet of the two lines, as can be checked by taking cofactor matrices. We call this double point the point of contact. Therefore, the polarity property expressed by Eq. (7) can be stated as
@@ 3.2 Quadrics and projective space
Similar definitions hold in projective 3-space $\mathbb{R}P^{3}$, using a 4-dimensional vector space $\mathtt{V}$ where now the dual space ${\mathbb{R}P^{3}}^{*}$ is the space of planes. A correlation is defined by the property that 3 collinear points map to 3 coaxial planes (three planes going through the same line: we use the word "axis" to emphasize the dual role that a line plays as the intersection of two planes in contrast to its more familiar role of connecting two points, for which we use the term “spear”. A polarity is defined the same way, but we now refer to the surface of points incident to their polar planes as a point-wise quadric. There are now 3 different cases to consider:
@@
The set of geometric elements represented by vectors in the kernel of $A$ is known as the degenerate set. For point-wise quadrics this is: the vertex of a cone, the points of the axis of a plane pair, the points of a double plane; and for plane-wise quadrics: the plane of a conic, the planes of the joining line of a point pair, the planes of a double point. For degenerate point-wise quadrics, the polarity mapping is only defined for points not in the degenerate set to planes, and vice versa for plane-wise quadrics. For point-wise quadrics:
@@ 3.2.2 Some contact lemmas
We conclude this section with a few lemmas that will be useful to prove the 8-quadric theorem. By ring contact triangle we mean a set of 3 ring contact pencils that pairwise have one conic in common.
@@
where $\alpha$ is a scalar (note that we are not entitled to arbitrarily scale it to $1$ because we already fixed the relative scaling of $A$ and $C$).
@@ 3.3 Notation
Before we turn to the proof of the theorems, we introduce some notation used in the sequel. First, an alternative notation for degenerate conics, used in Sec. 4.2 and Sec. 6. Secondly, notation for so-called complete quadrics, required to represent non-regular conics and quadrics, and used in Sec. 14.
@@ 3.3.2 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, as in Sec. 14, we employ the full notation introduced here.
@@ 4.1 Notation and terminology
Next, we introduce notation for the vertices and faces of the cube. Consult Figure 5, left. We decompose the cube vertices into two interlocked tetrahedra and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $\mathcal{F}^{ij}_{km}$ where $ijkm$ is a permutation of $0123$. The common point of the chords of contact in this face we write as $P^{ij}_{km}$, the face point of the face. Since the edges of the cube are double contact pencils, each edge contains a unique double line $\mathbf{{V}}^{i}_{j}$, the chord of contact of $\mathbf{{T}}^{i}$ and $\mathbf{{S}}_{j}$.
@@ 4.3 Penrose’s 8-conic theorem
Proof of Penrose-7 conic theorem.
@@ 4.4 Penrose’s 8-quadric theorem
Let $i$ and $j\in\{0,1,2,3\}$, $m$ and $n$ in $\{1,2,3\}$ and $P^{m}_{i}$ denote the ring point for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$, $p^{m}_{i}$ denote the corresponding ring plane, $\Lambda^{mn}_{ij}$ the common axis of the ring planes for the face $F^{mn}_{ij}$ and $\lambda^{mn}_{ij}$ the common spear (point-wise line) of the ring points. We call these the face axes, resp., the face spears. By assumption, they are well defined. Equivalently, given any one of the three complete faces of the 7-configuration, its four ring planes are not all identical. Consider, for instance, the face $\mathcal{F}^{12}_{03}$. By Lemma 5, we immediately conclude that $\mathbf{{S}}_{0}$ is not in ring contact with $\mathbf{{S}}_{3}$ and $\mathbf{{T}}^{1}$ is not in ring contact with $\mathbf{{T}}^{2}$. By similar reasoning applied to the other two complete faces, we get the following lemma:
@@
Since a face axis is determined by any two ring planes on a face, and every face of the 7-configuration has at least two ring planes, the axes and spears of the incomplete faces are also uniquely determined. In particular, the shared axis of the ring planes for the incomplete face $F_{ij}^{0n}$ must go through the ring planes for the two edges already given, so we set
@@ 4.4.4 Proof of Extrusion Lemma
Because of condition (ii) of the hypothesis of the 8-conic theorem, all the chords of contact for the face $(\mathbf{{S}}_{0},\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{S}}_{3})$ share a point whose representative in the basis $(P_{1},P_{2},P_{3})$ is necessarily in the common kernel of $V^{1}_{3}$, $V^{2}_{0}$ ,$V^{1}_{0}$ and $V^{2}_{3}$. However, by the regularity assumption on $\mathbf{{T}}^{2}$, $C$ has trivial kernel which is only consistent with Eq. (40) in the case $r=1$.
@@ 5 Examples of 8-quadric configurations
In the next example, we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus, each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the quadric $\mathbf{{S}}_{0}$ as an imaginary degenerate conic represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
@@ 7 Determinant structure and generalizations
Take $\Omega\subset\{1,2,...,n\}$. To proceed, we give names to some special subdeterminants (note: this is not exhaustive, an exhaustive description of the case $n=3$ will be given in the following section):

### v8

@@ 3.1 Conics and the projective plane
where the superscript T denotes the transpose, therefore these two associated line-point correlations coincide exactly when the point-line correlation is represented by a symmetric matrix $A=A^{T}$ which is known as a polarity. A polarity can also be characterized as a correlation that is also an involution. A line and point which are mapped to each other by a polarity are referred to as polar to each other.
@@ 3.1.1 Pencils of conics
Assume $A$ is regular. The pencil, with the exception of $A$, can be written $A+tB,t\in\mathbb{R}$. $A+tB$ is a degenerate conic exactly when $|A+tB|=0$, where vertical bars around a matrix denote the determinant. The left-hand side is a cubic polynomial in $t$, hence has at most 3 roots. ∎
@@ 3.1.2 Double contact
when $P$ is incident to $p$. In particular, the intersection points of $B$ with $A$ are always the intersection points of $p$ with $A$ and the tangent lines there are always the same, independently of $c$. Thus, $A$ and $B$ are said to be in double contact with the chord of contact $p$, and the pencil of conics defined by Eq. (6) is called a double contact pencil. The property of being in double contact is clearly self-dual. Indeed, two point-wise conics share tangent lines at two common points if and only if the two corresponding line-wise conics share turning points at two common lines. Furthermore, Eq. (6) is self-dual, that is, the line-wise conics of a double contact pencil can be expressed as the linear span of a regular conic and a double point, given by the meet of the two lines, as can be checked by taking cofactor matrices. We call this double point the point of contact. Therefore, the polarity property expressed by Eq. (7) can be stated as
@@ 3.2 Quadrics and projective space
Similar definitions hold in projective 3-space $\mathbb{R}P^{3}$, using a 4-dimensional vector space $\mathtt{V}$ where now the dual space ${\mathbb{R}P^{3}}^{*}$ is the space of planes. A correlation is defined by the property that 3 collinear points map to 3 coaxial planes (three planes going through the same line: we use the word "axis" to emphasize the dual role that a line plays as the intersection of two planes in contrast to its more familiar role of connecting two points, for which we use the term “spear”). A polarity is defined the same way, but we now refer to the surface of points incident to their polar planes as a point-wise quadric. There are now 3 different cases to consider:
@@
The set of geometric elements represented by vectors in the kernel of $A$ is known as the degenerate set. For point-wise quadrics this is: the vertex of a cone, the points of the axis of a plane pair, the points of a double plane; and for plane-wise quadrics: the plane of a conic, the planes of the joining line of a point pair, the planes of a double point. For degenerate point-wise quadrics, the polarity mapping is only defined for points not in the degenerate set to planes that go through the degenerate set, and vice versa for plane-wise quadrics. For point-wise quadrics:
@@ 3.2.2 Some contact lemmas
We provide a few lemmas that will be useful in the proof of the 8-quadric theorem. By ring contact triangle we mean a set of 3 ring contact pencils that pairwise have one conic in common.
@@
where $\alpha$ is a scalar (note that we are not entitled to arbitrarily scale it to $1$ because we already fixed the relative scaling of $A$ and $C$ with the first two equations).
@@ 3.3 Notation
Before we turn to the proof of the theorems, we introduce some notation used in the sequel. First, an alternative notation for degenerate conics, used in Sec. 4.2 and Sec. 6. Secondly, notation for so-called complete quadrics, required to represent non-regular conics and quadrics, and used in Theorem. 14.
@@ 3.3.2 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$ by equation (5), where by $\mathbf{{Q}}$ and $\mathbf{{Q}}^{*}$, we mean some matrices representing these quadrics. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, as in Sec. 14, we employ the full notation introduced here.
@@ 4.1 Notation and terminology
Next, we introduce notation for the vertices and faces of the cube. Consult Figure 5, left. We decompose the cube vertices into two interlocked tetrahedra and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $\mathcal{F}^{ij}_{km}$ where $ijkm$ is a permutation of $0123$. The common point of the chords of contact in this face we write as $P^{ij}_{km}$, the face point of the face. Since the edges of the cube represent double contact pencils, each such pencil contains a unique double line $\mathbf{{V}}^{i}_{j}=(\mathbf{p}^{i}_{j})^{2}$, where $\mathbf{p}^{i}_{j}$ the chord of contact of $\mathbf{{T}}^{i}$ and $\mathbf{{S}}_{j}$.
@@ 4.3 Penrose’s 8-conic theorem
Proof of Penrose-8 conic theorem.
@@ 4.4 Penrose’s 8-quadric theorem
Let $i$ and $j\in\{0,1,2,3\}$, $m$ and $n$ in $\{1,2,3\}$ and $P^{m}_{i}$ denote the ring point for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$, $p^{m}_{i}$ denote the corresponding ring plane, $\Lambda^{mn}_{ij}$ the common axis of the ring planes for the face $\mathcal{F}^{mn}_{ij}$ and $\lambda^{mn}_{ij}$ the common spear (point-wise line) of the ring points. We call these the face axes, resp., the face spears. By assumption, they are well defined. Equivalently, given any one of the three complete faces of the 7-configuration, its four ring planes are not all identical. Consider, for instance, the face $\mathcal{F}^{12}_{03}$. By Lemma 5, we immediately conclude that $\mathbf{{S}}_{0}$ is not in ring contact with $\mathbf{{S}}_{3}$ and $\mathbf{{T}}^{1}$ is not in ring contact with $\mathbf{{T}}^{2}$. By similar reasoning applied to the other two complete faces, we get the following lemma:
@@
Since a face axis is determined by any two ring planes on a face, and every face of the 7-configuration has at least two ring planes, the axes and spears of the incomplete faces are also uniquely determined. In particular, the shared axis of the ring planes for the incomplete face $\mathcal{F}_{ij}^{0n}$ must go through the ring planes for the two edges already given, so we set
@@ 4.4.4 Proof of Extrusion Lemma
Because of condition (ii) of the hypothesis of the 8-conic theorem, all the chords of contact for the face $\mathcal{F}^{12}_{03}$ share a point whose representative in the basis $(P_{1},P_{2},P_{3})$ is necessarily in the common kernel of $V^{1}_{3}$, $V^{2}_{0}$ ,$V^{1}_{0}$ and $V^{2}_{3}$. However, by the regularity assumption on $\mathbf{{T}}^{2}$, $C$ has trivial kernel which is only consistent with Eq. (40) in the case $r=1$.
@@ 5 Examples of 8-quadric configurations
In the next example, we consider three spheres $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ such that no sphere is completely contained in another. All spheres intersect the plane at infinity in the same imaginary circle $\mathbf{{S}}_{0}$, the absolute of Euclidean geometry. Thus, each sphere $\mathbf{{T}}^{i}$ is in ring contact with $\mathbf{{S}}_{0}$, where the conic $\mathbf{{S}}_{0}$ as an imaginary degenerate quadric represents at the same time the rings $\mathbf{{R}}_{0}^{i}$. We could proceed as in the previous example by choosing as quadrics $\mathbf{{S}}_{k}$ the intersection of the spheres $\mathbf{{T}}^{i}\cap\mathbf{{S}}_{0}$, which are circles (besides the absolute imaginary circle at infinity) and by slicing with a plane through the centers of the 3 spheres we would obtain again a Brianchon-like theorem, which turns out to be the theorem of the radical center of 3 circles.
@@ 7 Determinant structure and generalizations
Take $\Omega\subset\{1,2,...,n\}$. To proceed, we give names to some special subdeterminants (note: this is not exhaustive, an exhaustive description of the case $n=3$ will be given in Sec. 7.6):